# Path to Diversity: A Primer on ISAC-izing Commodity Wi-Fi for Practical Deployments

### v1

@@ V-E Summary: Towards Ubiquitous ISAC
This chapter has explored the dimension of Link Diversity, moving beyond single-link observations to exploit the macroscopic topology of the sensing network. We have navigated this landscape through two distinct paradigms:
  • Physical Link Diversity: By densifying the infrastructure with distributed transceivers, this approach physically expands the observation aperture. While it offers superior coverage and diversity (as seen in multi-static systems), it faces challenges in deployment cost and hardware synchronization.
  • Digital Link Diversity: By harnessing protocol-standardized feedback, i.e., BFI, this approach prioritizes ubiquity and accessibility. We highlighted the unique trade-off of BFI: while it sacrifices absolute ranging fidelity due to quantization, its inherent stability serves as a robust filter against hardware noise, making it promising for mobile sensing.
The case study of MUSE-Fi integrates the above concepts, demonstrating that practical ubiquity is not achieved by choosing one paradigm over the other, but by their synergy. By combining physical near-field isolation with digital sparse recovery, we can resolve complex multi-user interference on commodity devices.
Table VII summarizes these methodologies, highlighting their respective goals, enablers, and trade-offs. Having explored the potential of distributed multi-link topologies and ubiquitous digital feedback, our focus now shifts to the antenna domain. The following chapter investigates Spatial Diversity, examining how MIMO arrays and beamforming techniques enable directional sensing for ISAC.
@@ VI-B The Passive Paradigm: Rx-MIMO for Spatial Analysis
SpotFi [37], RoArray [16], Phaser [15], S-Phaser [19], M3 [8] High-Accuracy Localization
@@
BodyCompass [98], Wi-PIGR [101], WiWrite [47], AFall [4] Activity Recognition & User ID
@@ VI-B5 Spatial Representation Learning
  • Feature-Based Recognition: AFall [4] extracts explicit Spatial AoA profiles to distinguish falls via physics-based models. FallDeFi [52] aggregates time-frequency features from multiple antennas and classifies them via Support Vector Machines (SVM).
@@ VI-D Case Study: Realizing Programmable Directionality via Beamforming Feedback Control
Figure 17: The paradigm shift in MU-MIMO sensing. (a) Legacy systems struggle with the trade-off between throughput and sensing coverage. (b) The active paradigm (Beam-Fi) adopts a coordination strategy to simultaneously serve both by actively scheduling the beam direction [23]).
@@ VI-E Summary: Towards Directional ISAC
This chapter explored Spatial Diversity, the final degree of freedom determining angular resolution and energy distribution. Unlike Temporal and Frequency diversities which govern ranging, Spatial Diversity dictates the system’s ability to discriminate spatial geometry and, crucially, to actively steer electromagnetic resources. We mapped the evolution of this domain through two distinct paradigms:
  • Passive Rx-MIMO (Receiver-Side Analysis): Focusing purely on signal extraction, this paradigm employs advanced signal processing algorithms to resolve incoming angles from multipath reflections. While effective for tracking, it remains fundamentally constrained by passive blindness—it relies entirely on fortuitous illumination and is inherently vulnerable to energy nulls where targets become invisible.
  • Active Tx-MIMO (Programmable Directionality): Reversing the agency, this paradigm transforms the AP from a passive radiator into an active probe. By manipulating the feedback loop to steer beams, it turns interference management and power concentration into controllable sensing resources, ensuring that targets are actively highlighted rather than neglected.
Table IX summarizes these paradigms. As demonstrated by the Case Study, achieving true Directional ISAC requires transcending passive observation to embrace active control. By forging feedback to steer energy, we eliminate the blind spots of passive systems, realizing a high-fidelity sensing paradigm that coexists symbiotically with high-speed data transmission.
@@ VII-B Orchestration of Multi-Dimensional Diversity
Future architectures should explicitly address these systemic couplings, so that different diversity dimensions can reinforce rather than undermine one another.
@@ VII-C Resource Efficiency and Collaborative Coexistence
Collaborative Coexistence: As active sensing becomes prevalent, the network will face the severe mutual interference and spectral congestion. Uncoordinated active probing from multiple APs will lead to severe co-channel interference. We envision the development of collaborative ISAC protocols, where neighboring APs coordinate their sensing slots and spatial beams, akin to interference management in cellular networks, to ensure collaborative coexistence.
@@ VII-D Universal Generalization and Semantic Intelligence
A persistent bottleneck in Wi-Fi sensing is domain dependency—models trained in one environment often fail in another due to multipath variations [26]. While utilizing diversity mitigates this, a potentially transformative direction is to move beyond environment-specific training towards AI-native ISAC, which can be enabled by the following two promising approaches.
@@ VII-E The Privacy-Utility Paradox and Physical Security
As Wi-Fi evolves from a communication standard to a sensing modality, it inevitably raises new privacy challenges that must be addressed alongside its utility. In particular, these privacy challenges manifest in the following two aspects:
  • Plaintext Vulnerability: Standardized feedback features (e.g., BFI) are currently transmitted as plaintext management frames. This allows attackers to infer sensitive information without authentication, such as occupancy detection [85] or keystroke inference [27].
  • Active Sensing Risks: The transition to active modes transforms APs into programmable sensors. Without safeguards, attackers can exploit directional beams to “see” through walls, turning infrastructure into malicious surveillance tools.
Future work must develop physical layer privacy mechanisms, such as beamforming perturbation or friendly jamming. The goal is degrade sensing quality for unauthorized eavesdroppers while maintaining high fidelity for legitimate communication and authorized sensing applications.

### v2

@@ V-B4 Spatial Disentanglement (Multi-User Separation)
Paradigm Core Mechanism Representative Works Key Capability
Coverage & Triangulation Aggregates views to eliminate blind spots; exploits AoA, shadowing, and backscatter for geometric solving. WiSee [63], E-eyes [94], WiDraw [76], LiFS [90], WiDeo [44], [11] Robustness:
No blind spots.
View-Invariant Feature Extraction Combines multi-link Doppler to reconstruct absolute velocity vectors, or extracts BVP for domain independence. CARM [92], IndoTrack [48], WiTraj [100], RIM [15], WiDance [66], Widar3.0 [121], DeepMV [109] Generalization:
View/Location invariance.
3D Structure recovery.
Multi-user separation.
@@ V-E Summary: Towards Ubiquitous ISAC
This chapter explored Link Diversity, moving beyond single-link observations to realize Ubiquitous ISAC. We navigated this landscape through two paradigms:
  • Physical Link Diversity: Densifies infrastructure with distributed transceivers to physically expand the observation aperture. While offering superior coverage and diversity, it faces challenges regarding deployment cost and hardware synchronization.
  • Digital Link Diversity: Harnesses protocol-standardized feedback (BFI) to prioritize ubiquity. Despite sacrificing ranging fidelity due to quantization, BFI’s inherent stability effectively filters hardware noise, making it promising for mobile sensing.
The MUSE-Fi case study demonstrates that practical ubiquity relies on the synergy of these paradigms. By integrating physical near-field isolation with digital sparse recovery, the system resolves complex multi-user interference on commodity devices. Table VII summarizes these methodologies. Having established the potential of distributed topologies and digital feedback, we now shift to the antenna domain. The following chapter investigates Spatial Diversity, examining how MIMO arrays enable directional sensing for ISAC.
@@ VI-A3 The Challenge of Commodity Hardware
ArrayTrack [105] SpotFi [45], [16], RoArray [20], Phaser [19], S-Phaser [23], M3 [9] High-Accuracy Localization
@@
BodyCompass [115], Wi-PIGR [119], WiWrite [54], R-TTWD [122], AFall [5] Activity Recognition & User ID
VI-B The Passive Paradigm: Rx-MIMO for Spatial Analysis
@@ VI-B1 Stationary Physical Array Processing
  • The 2-D MP estimator [16] applies a search-free Matrix Pencil algorithm to jointly resolve TDOA and DOA via generalized eigenvalue decomposition.
@@ VI-B5 Spatial Representation Learning
  • Feature-Based Recognition: R-TTWD [122] applies PCA to extract statistical features for robust through-the-wall detection. AFall [5] extracts explicit Spatial AoA profiles to distinguish falls via physics-based models. FallDeFi [62] aggregates time-frequency features from multiple antennas and classifies them via Support Vector Machines (SVM).
@@ VI-D Case Study: Realizing Programmable Directionality via Beamforming Feedback Control
Figure 17: The paradigm shift in MU-MIMO sensing. (a) Legacy systems struggle with the trade-off between throughput and sensing coverage. (b) The active paradigm (Beam-Fi) adopts a coordination strategy to simultaneously serve both by actively scheduling the beam direction [29].
@@ VI-E Summary: Towards Directional ISAC
This chapter explored Spatial Diversity, the dimension governing angular resolution and energy steering. We mapped the evolution of this domain through two distinct paradigms:
  • Passive Rx-MIMO (Receiver-Side Analysis): Focuses on extracting spatial geometry via advanced signal processing. While effective for tracking, it is constrained by passive blindness—relying entirely on fortuitous illumination and remaining vulnerable to energy nulls.
  • Active Tx-MIMO (Programmable Directionality): Transforms the AP from a passive radiator into an active probe. By manipulating feedback to steer beams, it converts interference management into a controllable resource, ensuring targets are actively highlighted rather than neglected.
Table IX summarizes these paradigms. As demonstrated by the Case Study, achieving true Directional ISAC requires transcending passive observation to embrace active control, realizing a high-fidelity sensing paradigm that coexists symbiotically with high-speed data transmission.
@@ VII-B Orchestration of Multi-Dimensional Diversity
Future architectures should explicitly address these systemic couplings, so that different diversity dimensions can reinforce rather than undermine one another. Moreover, this orchestration can be extended beyond the transceiver to the environment itself. By integrating Reconfigurable Intelligent Surfaces (RIS) [118, 38, 35], future systems can artificially create additional link and spatial diversity, actively mitigating blockage and enhancing sensing resolution in complex scenarios.
@@ VII-C Resource Efficiency and Collaborative Coexistence
Collaborative Coexistence: As active sensing becomes prevalent, the network will face the severe mutual interference and spectral congestion. Uncoordinated active probing from multiple APs will lead to severe co-channel interference. We envision the development of collaborative ISAC protocols, where neighboring APs coordinate their sensing slots and spatial beams, akin to interference management in cellular networks, to ensure collaborative coexistence [24].
@@ VII-D Universal Generalization and Semantic Intelligence
While deep learning has transitioned Wi-Fi sensing from manual feature engineering to data-driven representation learning [61], a persistent bottleneck in Wi-Fi sensing is domain dependency—models trained in one environment often fail in another due to multipath variations [4, 37]. While utilizing diversity mitigates this, a potentially transformative direction is to move beyond environment-specific training towards AI-native ISAC, which can be enabled by the following two promising approaches.
@@ VII-E The Privacy-Utility Paradox and Physical Security
As Wi-Fi evolves from a communication standard to a sensing modality, it inevitably raises new privacy challenges that must be addressed alongside its utility [18]. In particular, these privacy challenges manifest in the following two aspects:
  • Plaintext Vulnerability: Standardized feedback features (e.g., BFI) are currently transmitted as plaintext management frames. This allows attackers to infer sensitive information without authentication, such as occupancy detection [102] or keystroke inference [2, 34].
  • Active Sensing Risks: The transition to active modes transforms APs into programmable sensors. Without safeguards, attackers can exploit directional beams to “see” through walls [89], turning infrastructure into malicious surveillance tools.
Future work must develop physical layer privacy mechanisms, such as beamforming perturbation or friendly jamming [33]. The goal is degrade sensing quality for unauthorized eavesdroppers while maintaining high fidelity for legitimate communication and authorized sensing applications.